OpenStudy (anonymous):

Find all solutions of the equation in the interval [0, 2π) algebraically. sin^2 x + cos x + 1 = 0

OpenStudy (anonymous):

\[ \sin ^2 x=1-\cos ^2 x \]

OpenStudy (anonymous):

Let \(u=\cos x\)

OpenStudy (anonymous):

what next...?

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

I don't know how to do it.

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

i got 0,π/2, k/π. Is that correct?!?

OpenStudy (anonymous):

@ybarrap Do you understand how to solve this my friend?

OpenStudy (ybarrap):

What happens when \(x=\pi\)?

OpenStudy (anonymous):

-1?

OpenStudy (ybarrap):

$$ \sin^2 \pi + \cos \pi + 1 = 0\\ \sin \pi = ?\\ \sin^2\pi=?\\ \cos \pi =?\\ $$

OpenStudy (anonymous):

sinπ= 0

OpenStudy (anonymous):

sin2π= 0 cosπ=-1

OpenStudy (ybarrap):

So is \(x=\pi\) the solution? 0-1+1=?

OpenStudy (anonymous):

1. got it. thanks